IF IAI=2,IBI=3 AND A(VECTOR)-B(VECTOR)=O,THEN-
(Ax(Ax(xA(AxB))))=?
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 

Grade:Select Grade

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear student,

Thanks for your post.

Let vector A is denoted by ‘A’ and vector b is denoted by ‘B’
Given, A – B = 0

Taking cross product with A on both sides of the equation,
A X (A – B) = A X 0
\RightarrowA X A – A X B = A X 0
\Rightarrow– A X B = 0
\RightarrowA X B = 0

Since, A X B = 0
\Rightarrow(Ax(Ax(xA(xAxB)))) = 0